OpenStudy (anonymous):

Solve 3cos^2θ + 5cosθ + 2 = 0 for 0° ≤ θ < 180°. ans:Ø,{109°},{132°},{180°}

OpenStudy (anonymous):

3cos^2θ + 5cosθ + 2 3(2cos^2(θ)-1)+5cosθ+2=0 6cos^2(θ)+5cosθ-1=0 (6cosθ-1)(cosθ+1)=0 cosθ=1/6 or cosθ=-1 since cosθ<180 we can neglect cosθ=-1 then right the general solution for cosθ=1/6.. You will get the answers :)